# Literature Filtering for Systematic Reviews with Transformers

### v1

@@ 4. Results
Table 5. Bottom 50% Accuracy by Model and Research Question
@@
Looking at the results in tables 4 and 5, the best-performing models tend to be toward the bottom of the table where we the modelling approaches make use of the pre-trained transformer models. There does not appear to be a single modelling approach that brings us toward optimal performance across all research questions, however, the fine-tuning approach resulted in improved performance across most research questions. Furthermore, this approach was the only approach to deliver functional AUC for the LANB and VERT questions. All other approaches provided performance that was barely better than random for these questions.
@@
5. Discussion
We conducted a general exploration of machine learning approaches to filtering articles returned in a bibliographic database search to identify evidence for automated inclusion or exclusion. We have conducted these experiments using a rigorous Leave-One-Question-Out methodology, which allows us to understand the ability of each approach to perform when presented with truly novel research questions.
Our experiments provides substantial evidence that the use of pre-trained transformer models offers promise for improving systematic review efficiency through general-purpose machine learning filters. After fine tuning, these models can generalise to recognise articles that are inappropriate for inclusion in the systematic review of genuinely novel research questions. When considering the use case of filtering returned articles to exclude the bottom 50%, the highest accuracy was achieved in 5 of the 10 questions by fine tuning the BioBERT model for a single epoch of training. In the remaining 5 questions, we obtained the highest (or equal highest) accuracy using the pre-trained BERT models to generate similarity match scores between the research questions and article abstract.
@@ 5. Discussion
Using Microsoft Word, readability statistics were generated for the standard format questions. Except for the LANB and VERT questions, all questions had a Flesch-Kincaid Grade Level (FKGL)(Kincaid1975, ) above the average level of 15.87 for journal articles, as reported by Kandula and Zeng-Treitlar (2008)(Kandula2008, ). As such, the language used by researchers when drafting questions aligns with a journal article corpus. The low FKGL scores may account for the general poor performance of the NLP models, except for Fine-tuned BioBERT, when analysing the LANB and VERT questions.
@@
Table 6. Comparison of Research Question Properties
@@
  • Key to Symbols:
    B=Broad, F=Focused, N=Named, ~=Not Defined, KW=selected keywords, MS=multi synonyms,
    DS=descriptive sentence, MO=Multiple Outcomes, SC=Standard Care, EX=Exclusions, GL=Grade Level
6. Conclusion
The transformer machine learning architecture offers unique advantages for building natural language processing tools. We have found that pre-trained BERT models can be employed to generate text comparison features that outperform a variety of standard approaches for the majority of systematic review questions investigated. In addition, we observed that fine tuning one of these models often resulted in the best of all surveyed methods. However, we also found that there was no single superior approach for all research questions.
Comparison of the nature of the ten research questions in our study demonstrated that, generally, questions that had substantial detail in the natural language description of the research question performed better. The presence of explicit exclusions in the research question may also pose problems for machine learning systems. The exception to these rules was the EMR question set that had neither exclusions nor a small limited description. The EMR problem required using the larger BERT model to achieve strong performance. The EMR question statement was the second largest in the dataset and contained a complete PICO characterisation, suggesting that research questions with richer detail may require larger pre-trained models. Whether larger models, additional data or fine-tuning can transform the BERT models into consistently superior models for all use cases remains to be seen.

### v2

@@ 4 Results
Table 6: Bottom 50% Accuracy by Model and Research Question
@@
Table 7: Accuracy of Top 50 Predictions by Model and Research Question
experiment CARGEL CIDP EMR ESG IORT LANB LMTA PBRT VERT PID
1 - Linear Baseline 0.04 0.14 0.08 0.04 0.04 0.18 0.22 0.08 0.06 0.02
1 - Naive Bayes (GNB) 0.02 0.12 0.1 0.0 0.0 0.3 0.18 0.08 0.02 0.08
1 - Naive Bayes (CNB) 0.04 0.1 0.1 0.04 0.0 0.34 0.24 0.02 0.04 0.04
1 - SVM Classifier 0.0 0.26 0.06 0.06 0.0 0.26 0.12 0.14 0.02 0.02
1 - ExtaTrees 0.02 0.06 0.04 0.06 0.0 0.3 0.14 0.06 0.14 0.02
1 - TFIDF-ExtaTrees 0.0 0.04 0.0 0.04 0.0 0.28 0.18 0.0 0.16 0.0
1 - TFIDF-LGBM 0.0 0.0 0.0 0.0 0.0 0.34 0.14 0.0 0.42 0.0
2 - Linear-PICO 0.04 0.16 0.04 0.14 0.04 0.28 0.16 0.3 0.02 0.04
2 - ExtaTrees-PICO 0.02 0.02 0.06 0.02 0.0 0.3 0.14 0.06 0.14 0.02
2 - TFID-ExtaTrees-PICO 0.0 0.02 0.0 0.04 0.0 0.34 0.16 0.04 0.08 0.0
2 - TFID-LGBM-PICO 0.0 0.0 0.0 0.0 0.0 0.22 0.1 0.04 0.38 0.0
3 - LGBM-BioBERT 0.0 0.0 0.0 0.02 0.0 0.34 0.12 0.0 0.34 0.0
3 - LGBM-BlueBERT 0.0 0.0 0.0 0.0 0.0 0.3 0.1 0.02 0.36 0.0
3 - LGBM-BlueBERTxL 0.0 0.0 0.0 0.02 0.0 0.3 0.12 0.0 0.3 0.0
3 - LGBM-BlueBERT-PICO 0.0 0.0 0.0 0.0 0.0 0.22 0.12 0.04 0.38 0.0
4 - Fine-tune-01e-BioBERT 0.0 0.12 0.04 0.0 0.0 0.02 0.08 0.1 0.02 0.1
4 - Fine-tune-02e-BioBERT 0.02 0.2 0.06 0.12 0.12 0.12 0.06 0.1 0.18 0.04
Looking at the results in tables 5 and 6, the best-performing models tend to be toward the bottom of the table where we the modelling approaches make use of the pre-trained transformer models. There does not appear to be a single modelling approach that brings us toward optimal performance across all research questions and by extension, all future research questions. However, the fine-tuning approach resulted in improved performance across most research questions. Furthermore, this approach was the only approach to deliver functional AUC for the LANB and VERT questions. All other approaches provided performance that was barely better than random for these questions.
@@
5 Discussion
This work has conducted a general exploration of machine learning approaches to filtering articles returned in a bibliographic database search to identify evidence for automated inclusion or exclusion. We have conducted these experiments using a rigorous Leave-One-Question-Out methodology, which allows us to understand the ability of each approach to perform when presented with truly novel research questions. Our observations from this work can be grouped into those associated with machine learning and those associated with the formation and phrasing of questions.
5.1 Machine Learning
Our work provides substantial evidence that the use of pre-trained transformer models offers promise for improving systematic review efficiency through general-purpose machine learning filters. After fine tuning, these models can generalise to recognise articles that are inappropriate for inclusion in the systematic review of genuinely novel research questions. When considering the use case of filtering returned articles to exclude the bottom 50%, the highest accuracy was achieved in 5 of the 10 questions by fine tuning the BioBERT model for a single epoch of training. In the remaining 5 questions, we obtained the highest (or equal highest) accuracy using the pre-trained BERT models to generate similarity match scores between the research questions and article abstract.
However, the ideal methodology for building a general-purpose filter is unclear. For some questions, additional rounds of fine-tuning improved performance, whereas for others, it did not. When the fine-tuned model did not provide optimal accuracy, we found that different BERT models (or methods such as using PICO text) resulted in optimal performance. It remains to be seen whether there is a set of parameters (training rounds and hyper-parameters) or a methodology that will result in an optimal general-purpose filter.
Our comparisons with standard tabular data machine-learning algorithms showed that the BERT-based modelling improved results when considering the AUC or accuracy in the bottom 50%. However, when we looked at the ability of the models to identify the most likely articles for inclusion, judged by accuracy in the top 50 predictions, we found that simple linear models outperformed all others. The best approach involved using a PICO representation of the research questions and then generating text match features with the PICO elements independently. This surprising result emphasises that the selection of machine learning approaches depends heavily on the model’s intended use.
We will continue to explore other approaches to fine-tuning these models, but we have also conducted an examination of what made certain research questions easier or harder to solve. These insights can help guide the development of tools that might satisfy certain subsets of all systematic review questions.
5.2 Research Question Phrasing
@@
For all standard questions, the Population and Intervention elements of the PICO format were included. Indicating that all questions had the two primary PICO elements that are used in the initial screen for article selection for full-text review.
Using Microsoft Word, readability statistics were generated for the standard format questions. Except for the LANB and VERT questions, all questions had a Flesch-Kincaid Grade Level (FKGL)[48] above the average level of 15.87 for journal articles, as reported by Kandula and Zeng-Treitlar (2008)[49]. As such, the language used by researchers when drafting questions aligns with a journal article corpus. The low FKGL scores may account for the general poor performance of the NLP models, except for Fine-tuned BioBERT, when analysing the LANB and VERT questions.
@@
Within the PICO elements, the Population, Intervention and Comparator structure vary across research questions. These variations reflect differences in research topics, the health technology being investigated, and the comparator technology considered to be the current clinical practice. In contrast, the structure of the outcome parameter adopted by researchers was a descriptive sentence that included a list of the relevant outcomes. For the PBRT and PID questions, the PICO format in experiment 2 returned the highest AUC (Table 5). As such, there appears to be some benefit of having separate statements for the PICO elements over the standard question format in these scenarios. This is emphasised by the fact that in experiment 4 these same questions performed very poorly with a fine-tuned BERT model.
@@
Given the importance of the Population and Intervention in the initial screening process, experiments are needed to define the optimal format, level of content and FKGL for the population and Intervention elements in a simple natural language question or the individual PI categories.
Table 8: Comparison of Research Question Properties
@@
Notes: B=Broad, F=Focused, N=Named, ~=Not Defined, KW=selected keywords, MS=multi synonyms, DS=descriptive sentence, MO=Multiple Outcomes, SC=Standard Care, EX=Exclusions, GL=Grade Level
6 Conclusion
The transformer machine learning architecture offers unique advantages for building natural language processing tools. We have found that pre-trained models can be employed to generate text comparison features that outperform a variety of standard approaches for the majority systematic review questions investigated. In addition, we observed that fine tuning one of the models for the specific task of identifying suitable articles for a given research question could result in the best of all surveyed method. However, we also found that there was no single superior approach for all research questions. The natural variety in both the phrasing and specificity of these research questions prevented determination of a single ideal modelling approach that could be expected to perform optimally for all research questions.
Comparison of the nature of the ten research questions in our study demonstrated that, generally, questions that had substantial detail in the natural language description of the research question performed better. The presence of explicit exclusions in the research question may also pose problems for machine learning systems. The exception to these rules was the EMR question set that had neither exclusions nor a small limited description. The EMR problem required using the larger BERT model to achieve strong performance. The EMR question statement was the second largest in the dataset and contained a complete PICO characterisation, suggesting that research questions with richer detail may require larger pre-trained models.
We have demonstrated that for some research questions using the PICO framework is able to deliver superior performance over simple natural language expressions. This suggests that in some scenarios, the machine learning models benefit from pre-categorising content before generating the classification model’s natural language features. This result suggests that custom Transformer architectures (allowing multiple independent text inputs) might provide superior performance.
Overall, judged by AUC, the BERT-based models are consistently superior. When considering a usage process like the automatic exclusion of articles, we look at the accuracy in the bottom 50% metric. By this account, the BERT models are still generally the strongest performer, although, for certain research questions, they are outperformed by tree-based learning algorithms. However, if we change the usage pattern to consider automatic inclusion of the top 50 predictions, then BERT models perform poorly. For this use case, simple linear models provide the best differentiation. Whether additional data or fine-tuning can transform the BERT models into consistently superior models for all use cases remains to be seen.